ka for ch3cooh is 1.8 find out the percentage dissociation of o.2 m ch3cooh in 0.1 m hcl

Grade:12th pass

1 Answers

Arun
4 years ago
HCl dissociated completely. 

Now, let the concentration of acetate ion be x. Therefore, hydrogen ion contribution from acetic acid = x.

Since HCl dissociates completely, hydrogen ion contribution from HCl =0.1M

Total [H+] concentration = x + 0.1

Now, Ka for acetic acid = 1.8 x 10-5
 
1.8 x 10-5  = x (x + 0.1)/0.2
on solving x = 0.35 * 10^-4
 
hence percentage dissociation = 0.35 * 10^-4 * 100 /0.2 = 0.0175  
Hope it helps. Let us know if there is any issue.